A Construction company XYZ purchased a Crane for $600,000 and it will have a Salvage value of $60,000 at the end of its useful life. Company XYZ uses a 150% Declining-Balance Depreciation Method with a Switchover to Straight-Line Depreciation Method. Calculate the required Depreciation amounts for every year, identify the year that the Switchover will occur to Straight-Line Method, and fill all required amounts in the given TABLE below. List all your calculations below the table.

Year, K

Please show all calculations, and please no Excel usage. Will rate. Thank you.
Answer #1
Year Book Value Depreciation 150% SLM Depreciation BV EoY
0 $ 600,000 - - - $ 600,000
1 600,000 0.1875×600,000 =112,500 0.125*(600000-60000) = 67,500 112,500 487,500
2 487,500 0.1875*487,500 = 91406.25 (487500-60000)/7= 61,071.43 91,406.25 396,093.75
3 396,093.75 74,267.58 (396,093.75-60000)/6=56,015.625 74,267.58 321,826.17
4 321,826.17 60,342.41 (321,826.17-60,000)/5= 52,365.23 60,342.41 261,483.46
5 261,483.46 49,028.21 (261,483.46-60000)/4)= 50,370.94 50,370.94 211,112.82
6 211,112.82 39,583.65 50,370.94 50,379.94 160,741.88
7 160,741.88 30,139.10 50,370.94 50,370.94 10,370.94
8 110,370.94 20,694.55 50,370.94 50,370.94 60,000

Straight line method depreciation charges = (1/8) = 12.5% every year

150% DB depreciation = 1.5 × 12.5 = 18.75% every year

Here useful life is not given thus I have assumed it by counting number of rows provided. PleaSe contact if having any doubt.

Please refer the attached picture solved it using excel too got the same answer

0 600000 1 600000 112500 67500 112500 487500 2 487500 91406.25 61071.43 91406.25 396093.75 74267.58 3 396093.75 56015.63 7426

Please contact if having any query will be obliged to you for your generous support. Please help me it mean a lot to me. Thank you.